# Enhancing Multistep Brent Oil Price Forecasting with a Multi-Aspect Metaheuristic Optimization Approach and Ensemble Deep Learning Models

### v1

@@ Abstract
Accurate crude oil price forecasting is crucial for various economic activities, including energy trading, risk management, and investment planning. Although deep learning models have emerged as powerful tools for crude oil price forecasting, achieving accurate forecasts remains challenging. Deep learning models’ performance is heavily influenced by hyperparameters tuning, and they are expected to perform differently under various circumstances. Furthermore, price volatility is also sensitive to external factors such as world events. To address these limitations, we propose a hybrid approach combining metaheuristic optimisation and an ensemble of five popular neural network architectures used in time series forecasting. Unlike existing methods that apply metaheuristics to optimise hyperparameters within the neural network architecture, we exploit the GWO metaheuristic optimiser at four levels: feature selection, data preparation, model training, and forecast blending. The proposed approach has been evaluated for forecasting three-ahead days using real-world Brent crude oil price data, and the obtained results demonstrate that the proposed approach improves the forecasting performance measured using various benchmarks, achieving 0.000127 of MSE.
@@
Crude oil price forecasting , Brent oil forecasting , Grey Wolf Optimizer , ensemble model , time series forecasting.
journal: Applied Energy
[…]
[uinvr]organization=University of Verona,country=Italy
@@ 1 Introduction
Time series forecasting plays a crucial role in various industries, from finance to logistics to energy, as it provides valuable insights into future trends and helps businesses make informed decisions. However, crude oil forecasting is particularly challenging due to the complex dynamics of the oil market and the non-linear and volatile nature of crude oil prices.
In recent years, deep learning techniques have gained prominence for their ability to capture intricate patterns in time series data. Long Short-Term Memory (LSTM) and Gated Recurrent Unit (GRU) networks have emerged as popular choices for modelling temporal dependencies in such forecasting tasks [1, 2]. Their success is attributed to their inherent ability to handle long-range dependencies, making them particularly suitable for time series data with varying temporal patterns.
Nonetheless, achieving accurate predictions remains a challenge. Individual deep-learning models typically struggle to capture all data patterns, and their performance is heavily dependent on the training configuration. The effectiveness of individual models varies depending on the given dataset and conditions. Furthermore, many external factors influence crude oil prices, including geopolitical events and global economic conditions [3, 4, 5, 6].
In other words, the capacity to enhance the capture of dataset patterns, identify external factors influencing the formation of these patterns, and select the appropriate network architecture while configuring it optimally are critical factors that significantly impact forecasting performance. The optimal choice often varies depending on the dataset’s characteristics and the specific forecasting problem at hand.
To address these challenges, we propose a hybrid approach to combine the Grey Wolf Optimizer (GWO) with an ensemble of five deep learning networks- LSTM, GRU, CNN-LSTM, CNN-LSTM-attention, and an encoder-decoder-LSTM. GWO is a bio-inspired metaheuristic optimisation algorithm which simulates the leadership hierarchy and grey wolf hunting mechanism in nature. The objective is to make a significant contribution to the advancement of research in the field of crude oil forecasting, building upon the recognition of the challenges mentioned above and the advantages of meta-heuristic optimisation and conditional forecast combination. This approach can be particularly beneficial in forecasting tasks where different models of diverse structures and data assumptions are expected to perform differently under various conditions. For example, one neural network can perform better than another in a specific dataset and vice-versa.
@@
Metaheuristic optimisation techniques present a powerful solution for optimal hyperparameters tuning [7, 8]. Metaheuristics excels in reaching near-optimal solutions with limited information and computational resources. These algorithms generally give better results and converge faster than blind approaches like Exhaustive Grid Search. This makes them particularly valuable in optimising deep learning methods for time series forecasting.
Although some researchers have already exploited metaheuristic optimisation for tunning hyperparameters (such as epochs numbers, batch size) of ML models to improve WTI crude oil forecasting, i.e. [9], we believe this exploitation can be extended to optimise the entire process including parameters outside the ML network structures such as feature selection, defining the suitable sliding window, and to fuse multiple forecasts.
@@
2 Related works
Methods used to forecast crude oil prices can be categorised into three primary groups: statistical, econometric, and machine learning-based approaches. Statistical methods provide a foundation for understanding and analysing historical oil price data. They are handy for identifying trends, seasonal patterns, and cyclical fluctuations in oil prices [10, 11]. Common statistical techniques employed in crude oil forecasting include time series analysis and statistical forecasting models, such as autoregressive integrated moving average (ARIMA) [11]. Statistical methods are relatively simple to understand and implement. However, statistical methods are limited in capturing the nonlinear and unpredictable nature of oil price movements, especially in response to exogenous events. Econometric methods are grounded in statistical theory and economic principles, and they are particularly useful for understanding the causal relationships between economic variables and oil prices. Vector Autoregression (VAR) is a common econometric model used in crude oil forecasting [12, 13]. However, Econometric models often make assumptions about the underlying data distribution and the relationship between variables. Violations of these assumptions lead to less reliable forecasts.
Deep learning techniques have become increasingly popular in the field of time series forecasting in general due to their ability to capture non-linear and more complex patterns in the data [14, 15]. Researchers have harnessed various deep learning architectures, including long short-term memory (LSTM) networks [1, 2, 16, 17], gated recurrent units (GRUs) [18, 17, 19], convolutional neural networks (CNNs) [20, 21], and transformer models [22], to forecast crude oil prices accurately.
In recent years, GRU (Gated Recurrent Unit) and LSTM (Long Short-Term Memory) networks have emerged as prominent choices for deep learning applications in crude oil price forecasting due to their ability to capture long-term dependencies and handle sequential data effectively. Several studies have demonstrated the superiority of GRU and LSTM networks for crude oil price forecasting [14]. To further enhance their forecasting capabilities, researchers have explored various hybrid and combination methods, such as integrating convolutional neural networks (CNN) and attention mechanisms into GRU and LSTM architectures [20, 19, 18, 23, 24, 25, 26]. Additionally, incorporating external factors has been a focus, with studies demonstrating the effectiveness of these strategies in refining crude oil price predictions [27, 28, 25, 29].
However, the efficacy of these approaches often hinges on the quality and comprehensiveness of feature engineering and the availability of extensive historical data [27], in addition to the optimal hyperparameters setup. Recently, metaheuristic algorithms have been popular for optimising machine learning models training [7, 30]. These algorithms have been employed to some extent for optimising ML models for crude oil forecasting, i.e. [9], leading to improved prediction accuracy. On the other hand, This work falls within the few research works that utilise the Brent crude oil database, as most research typically relies on the Texas crude oil database.
@@ 3 Preliminaries
3.1 Metaheuristic optimisation
@@ 3.2 Grey Wolf Optimizer (GWO)
The Grey Wolf Optimizer (GWO) is a bio-inspired swarm intelligence optimisation algorithm inspired by grey wolves, which has been introduced in [35] in 2014. This algorithm mimics the leadership hierarchy and hunting mechanism of grey wolves in nature. To simulate the leadership hierarchy in the GWO, four types of grey wolves—alpha $(\alpha)$, beta $(\beta)$, delta $(\delta)$, and omega $(\omega)$—are employed. These categories simulate the leadership hierarchy, where the optimisation process is guided by three wolves ($\alpha,\beta$, and $\delta$), while the omega $(\omega)$ wolves follow. As observed in the hunting process, grey wolves tend to encircle their prey, a phenomenon mathematically modelled by equations (1) and (2).
@@
3.3 Deep learning architectures for time series forecasting
LSTM (Long Short-Term Memory) and GRU (Gated Recurrent Unit) are two types of recurrent neural networks (RNNs) that have gained immense popularity in time series forecasting due to their ability to capture temporal dependencies in data effectively. These dependencies, often characterised by patterns that persist over time, are crucial for accurate forecasting. LSTM and GRU have demonstrated superior performance in various time series forecasting tasks, ranging from energy price prediction to weather forecasting [14]. Their ability to handle complex temporal dynamics and relatively simple architecture compared to more advanced models like transformers and GANs have made them the go-to choice for many time series forecasting applications. Researchers are constantly exploring ways to enhance the forecasting accuracy of LSTM and GRU models by combining them with convolutional neural networks (CNNs) to extract spatial features from time series data, incorporating attention mechanisms to focus on relevant portions of the input sequence, incorporating relevant external factors, and integrating statistical approaches to leverage the strengths of both neural networks and traditional forecasting methods [20, 23, 24, 25, 27, 26]. In this work, we examine the following popular deep-learning architectures:
  • 1.
@@ 3.3 Deep learning architectures for time series forecasting
To sum up, We train five models to predict three ahead-days, using the GWO metaheuristic optimisation to optimise training hyperparameters and incorporating two external factors (USDX and SENT) as described in the dataset section.
4 Proposed approach
In this work, We propose to combine the Grey Wolf Optimizer (GWO) with an ensemble of five deep learning networks- LSTM, GRU, CNN-LSTM, CNN-LSTM-attention, an encoder-decoder-LSTM for enhancing Brent crude oil price forecasting.
Figure 2: Proposed Approach
The aim is to enhance the predictive accuracy of crude oil prices by leveraging the acknowledged benefits of both metaheuristic optimisation and ensemble forecasting methods. The proposed approach is implemented in two main stages, as shown in the figure (2). In the first stage, we train five deep neural networks, proven for their proficiency in handling time series prediction. In this stage, the GWO algorithm is used to find optimal parameters for tuning the neural networks’ training. This stage also encompasses optimisation of feature selection and determining the window size. In the second stage, The GWO is utilised to blend the predictions obtained in the first stage at the optimal configuration for each neural network. To implement our approach, we build our code upon mealpy library [34], where users can select from a wide range of advanced population-based metaheuristic algorithms.
4.1 Multi-aspect metaheuristic optimisation
The Grey Wolf Optimization (GWO) algorithm is employed at four different stages to optimise crucial aspects of the ensemble forecasting process: feature selection, sliding window size determination, model hyperparameter tuning, and final forecast blending. This systematic approach ensures that the ensemble model leverages the strengths of all constituent models and produces highly accurate predictions
@@ 4.1.1 Feature selection
Forecasting crude oil prices is significantly influenced by external factors, which play a pivotal role in shaping the market dynamics. Factors such as geopolitical events, economic indicators, global demand-supply imbalances, and environmental policies can substantially impact oil prices [29]. Understanding and incorporating these external factors into forecasting models is crucial for accurate predictions, as they contribute to the volatility and unpredictability of the oil market. USD index rate and sentiment analysis index are two popular external variables that have been widely applied by researchers in the tasks of forecasting crude oil price [6, 26, 4, 38]. The USDX index, reflecting the performance of the U.S. dollar against a basket of major currencies, offers valuable information on currency strength and global economic conditions. Considering the substantial influence of currency fluctuations on oil prices, the USDX index is a crucial indicator. Simultaneously, sentiment analysis, which gauges market sentiment from news and social media, contributes to understanding market psychology and potential price movements. Integrating eternal factors enhances the model’s ability to capture complex relationships and dynamic market influences, consequently improving the accuracy of crude oil price forecasts [27, 28].In this work, we examined integrating both indices, USDX and sentiment score, individually and combined.
4.1.2 Determination of sliding window size
The determination of the sliding window size plays a pivotal role in time series forecasting. This parameter represents the temporal scope of historical data used to train a model, influencing its ability to capture patterns and trends. Selecting an optimal sliding window size is a delicate balance between incorporating sufficient historical information for meaningful pattern recognition and avoiding overfitting or loss of relevance. A smaller window may lead to overlooking essential patterns, while a more significant window might introduce noise or outdated information. Thus, carefully considering and optimising the sliding window size is crucial to ensure the trained model learns from past data effectively, contributing to more accurate and reliable crude oil price forecasts.
4.1.3 Tuning of model hyperparameters
Hyperparameter tuning, the process of carefully adjusting the settings of a deep learning model, is crucial for optimal performance of deep learning networks. Finding optimal hyperparameters (such as learning rate, optimiser, hidden and unite number) significantly impacts training efficiency and model accuracy. However, tuning these hyperparameters is challenging due to their effects’ interplay and the search space’s high dimensionality. Manual exploration can be time-consuming and error-prone, while automated techniques like grid search and Bayesian optimisation can be computationally expensive. The emergence of metaheuristic algorithms like genetic algorithms and particle swarm optimisation offers promising avenues for efficient hyperparameter tuning. These algorithms can explore the search space more effectively and efficiently, leading to better-performing models.
@@ 4.1.4 Blending of final forecasts
Ensemble forecast methods can enhance accuracy by combining forecasts of diverse models that capture different aspects of the data. This often leads to better generalisation and robustness. However, the success of the ensemble methods depends on the individual models’ diversity and quality and the nature of the data. In practice, ensemble methods are known to provide performance improvements in many cases, but there are situations where individual models might perform equally well or even outperform an ensemble. Variant-Weight technique is a well-established method for ensemble forecasting, in which multiple models are combined by assigning distinct weights to their respective predictions [24, 23]. The success of this method hinges on identifying the optimal weight allocation for each model. To this end, we employ GWO to optimise the weight distribution for our target models.
@@ 5 Experiments
5.1 Dataset
The dataset used for the experiments consists of the daily observations of the Brent crude oil price for the period from (January 3, 2012, to April 1, 2021), obtained from Bloomberg, including the period where the COVID-19 pandemic has affected energy and stock markets. It also includes Two external factors: the daily USD closing price (USDX) from Bloomberg and a daily cumulative sentimental score (SENT) obtained by the CrudeBERT model described in [39]. CrudeBERT is a variant of FinBERT that has been fine-tuned towards assessing the impact of market events on crude oil prices, focusing on frequently occurring market events and their impact on market prices according to Adam Smith’s theory of supply and demand. Mainly, CrudeBERT dataset used headlines originating from 1034 unique news sources, published on the Dow Jones newswires (approx. 21,200), Reuters (approx. 3,000), Bloomberg (approx. 1,100), and Platts (approx. 870). The sentimental score data is described and publicly available in [39]. The period from 2012 to 2021 was chosen for analysis due to the availability of sentiment scores during this time frame. Figure 3 displays the time-series plot of the daily closing price of Brent for this period. The dataset has been cleaned, normalised, and then split as follows: 80% for training, 10% for validation, and 10% for testing.

### v2

# Multistep Brent Oil Price Forecasting with a Multi-Aspect Meta-heuristic Optimization and Ensemble Deep Learning Model

@@ Abstract
Accurate crude oil price forecasting is crucial for various economic activities, including energy trading, risk management, and investment planning. Although deep learning models have emerged as powerful tools for crude oil price forecasting, achieving accurate forecasts remains challenging. Deep learning models’ performance is heavily influenced by hyperparameters tuning, and they are expected to perform differently under various circumstances. Furthermore, price volatility is also sensitive to external factors such as world events. To address these limitations, we propose a hybrid approach that integrates metaheuristic optimization with an ensemble of five widely used neural network architectures for time series forecasting. Unlike existing methods that apply metaheuristics to optimise hyperparameters within the neural network architecture, we exploit the GWO metaheuristic optimiser at four levels: feature selection, data preparation, model training, and forecast blending. The proposed approach has been evaluated for forecasting three-ahead days using real-world Brent crude oil price data, and the obtained results demonstrate that the proposed approach improves the forecasting performance measured using various benchmarks, achieving 0.000127 of MSE.
@@
Crude oil price forecasting, Brent oil analysis, time series forecasting, Ensemble learning, Grey Wolf Optimizer
@@ 1 Introduction
Time series forecasting plays a crucial role in various industries, from finance to logistics to energy, as it provides valuable insights into future trends and helps businesses make informed decisions. However, crude oil forecasting is particularly challenging due to the complex dynamics of the oil market and the non-linear and volatile nature of crude oil prices. In recent years, deep learning techniques have gained significant recognition for their ability to capture complex patterns in time series data. Long Short-Term Memory (LSTM) and Gated Recurrent Unit (GRU) networks have become popular choices for modeling temporal dependencies in forecasting tasks [1, 2]. Despite the emergence of more recent deep learning models such as transformers and generative models, LSTM and GRU remain popular choices for time series forecasting. LSTM and GRU networks excel in tasks with clear temporal sequences [3, 4]. Moreover, they are more accurate and quicker to train on small datasets, making them well-suited for daily frequency datasets such as daily oil prices [3, 5, 6]. To further enhance their performance for time series forecasting, researchers have increasingly integrated additional components, such as CNN-LSTM architectures, combining convolutional neural networks with LSTM to capture spatiotemporal patterns [7]. Nonetheless, achieving accurate predictions remains a challenge. Individual deep-learning models typically struggle to capture all data patterns, and their performance is heavily dependent on the training configuration. The effectiveness of individual models varies depending on the given dataset and conditions. Furthermore, many external factors influence crude oil prices, including geopolitical events and global economic conditions [8, 9, 10, 11]. In other words, the capacity to enhance the capture of dataset patterns, identify external factors influencing the formation of these patterns, and select the appropriate network architecture while configuring it optimally are critical factors that significantly impact forecasting performance. The optimal choice often varies depending on the dataset’s characteristics and the specific forecasting problem at hand. To address these challenges, we propose a hybrid approach to combine the Grey Wolf Optimizer (GWO) with an ensemble of five deep learning networks- LSTM, GRU, CNN-LSTM, CNN-LSTM-attention, and an encoder-decoder-LSTM. GWO is a bio-inspired metaheuristic optimisation algorithm which simulates the leadership hierarchy and grey wolf hunting mechanism in nature. The objective is to make a significant contribution to the advancement of research in the field of crude oil forecasting, building upon the recognition of the challenges mentioned above and the advantages of meta-heuristic optimisation and conditional forecast combination. This approach can be particularly beneficial in forecasting tasks where different models of diverse structures and data assumptions are expected to perform differently under various conditions. For example, one neural network can perform better than another in a specific dataset and vice-versa.
@@
Metaheuristic optimisation techniques present a powerful solution for optimal hyperparameters tuning [12, 13]. They excels in reaching near-optimal solutions with limited information and computational resources. These algorithms generally give better results and converge faster than blind approaches like Exhaustive Grid Search. This makes them particularly valuable in optimising deep learning methods for time series forecasting. While some researchers, such as [14], have already employed metaheuristic optimization algorithms for hyperparameter tuning in deep learning models for crude oil forecasting, we extend this application by utilizing the Grey Wolf Optimizer (GWO) algorithm to optimise the entire process including parameters outside the ML network structures such as feature selection, sliding window, and to fuse multiple forecasts.
@@
2 Litrature Review
Initially, statistical models such as autoregressive integrated moving average (ARIMA) [15] were applied to forecast oil prices. As statistical methods relies on mathematical assumptions, they struggle to capture the nonlinear and intricate dynamic of oil prices. Recent years have witnessed strong growth in adopting deep learning for time-series forecasting, driven by its remarkable ability to capture more complex patterns within the data. Researchers have employed various deep learning networks, including long short-term memory (LSTM) networks [1, 2, 16, 17], gated recurrent units (GRUs) [18, 17, 19], convolutional neural networks (CNNs) [20, 21], and transformer models [22], to forecast crude oil prices accurately based on historical price data and additional data sources such as sentiment analysis, supply and demand data [23, 24]. A popular approach combines various neural networks layers to improve crude oil forecasting. For instance, many researchers have merged LSTM networks with CNN and self-attention mechanisms to capture temporal, local, and long-term dependencies in historical price data [20, 19, 18, 25, 26, 27, 28]. Among these deep learning networks, GRU and LSTM stand out as the most widely applied methods for crude oil price forecasting. These methods excel in capturing temporal dependencies. Additionally, they are well-suited for smaller datasets and are computationally efficient compared to the more complex architectures of Transformers. Readers can refer to [6, 7, 4] for further details on this topic. [3] conducted an empirical analysis of various deep learning architectures, including LSTM, GRU, CNN-LSTM, and Transformer-based models, for oil price forecasting, demonstrating the superiority of GRU and LSTM. Since tuning model hyperparameters is crucial for enhancing prediction performance, metaheuristic algorithms have recently gained popularity for optimizing the training of machine learning models [12, 29]. These algorithms have been employed to some extent for optimising ML models for crude oil forecasting, i.e. [14], leading to improved prediction accuracy. In this work, we present a comprehensive approach to addressing the aforementioned limitations from multiple perspectives by developing an optimized ensemble model based on established models in the literature.
@@ 3 Preliminaries
3.1 Metaheuristic optimisation for machine learning models
@@ 3.2 Grey Wolf Optimizer (GWO)
The Grey Wolf Optimizer (GWO) is a bio-inspired swarm intelligence optimisation algorithm inspired by grey wolves, which has been introduced in [34] in 2014. This algorithm mimics the leadership hierarchy and hunting mechanism of grey wolves in nature as illustrated in Figure 1. To simulate the leadership hierarchy in the GWO, four types of grey wolves—alpha $(\alpha)$, beta $(\beta)$, delta $(\delta)$, and omega $(\omega)$—are employed. These categories simulate the leadership hierarchy, where the optimisation process is guided by three wolves ($\alpha,\beta$, and $\delta$), while the omega $(\omega)$ wolves follow. As observed in the hunting process, grey wolves tend to encircle their prey, a phenomenon mathematically modelled by equations (1) and (2).
@@
4 Proposed approach
In this work, we propose a deep learning ensemble forecasting approach, denoted as GWO-ensemble, to enhance multi-step Brent oil price forecasting. Since hyperparameter tuning is a crucial factor affecting model performance, we integrate the well-known Grey Wolf Optimizer (GWO). Unlike existing models that focus solely on neural network hyperparameter tuning, we apply GWO throughout all stages of model development, including feature selection, sequence data preparation, training individual models, and blending forecasts to generate the final output.
The proposed approach is implemented in two main stages, as shown in the figure 2. In the first stage, we train five deep neural networks (described in Section 4.1), proven for their proficiency in handling time series prediction. In this stage, GWO is used for optimisation of selecting sliding windows for building data sequences, feature selection, and tuning neural network hyperparameters.
The second stage involves blending the forecasts of the individual models using a weight-variant technique. In this stage, GWO is employed to optimize the weights for the forecast blending process.
Figure 2: Proposed Approach
To implement the proposed model, we build our code upon mealpy library [33], where users can select from a wide range of advanced population-based metaheuristic algorithms.
In Subsection 4.1, we describe the individual models used as the bases for the ensemble learning. Then, in the next subsection 4.2, we explain the ensemble learning approach. In Subsection 4.3, we outline aspects and stages in which the GWO optimiser has been employed.
4.1 Individual models
Our ensemble forecasting model is developed using blended ensemble learning that combines the following five established deep learning architectures. First, we trained each model individually, and then the individual forecasts obtained by each model were blended using a weighted-variant method, as illustrated in Figure 2.
  1. 1.
@@ 4.1 Individual models
4.2 GWO ensemble model
Inspired by the advantages of blending forecasts using conditional methods, we explore a weighted ensemble method combined with a meta-heuristic optimiser to fuse forecasts obtained from individual models. The weighted ensemble method in forecasting involves a mathematical approach to combine the predictions of multiple individual models to improve overall accuracy and robustness. Each model in the ensemble contributes to the final forecast with a certain weight. The final ensemble forecast is then obtained by summing the products of each model’s forecast with its corresponding optimal weight. Mathematically, this can be represented as the weighted sum:
Weighted ensemble prediction $=\sum_{i=1}^{n}M_{i}W_{i}$ (10)
Here, $W_{i}$ represents an optimal weight assigned to $model_{i}$, where $0\leq W_{i}\leq 1$, and $\sum W_{i}=1$. While $M_{i}$ represents $model_{i}$ forecasts. In this study, a dynamic weighted ensemble method was employed to blend the predictions of the top five forecasts obtained from the SENT-Bi-GRU, SENT-Bi-LSTM, SENT-CNN-Bi-LSTM, SENT-CNN-Bi-LSTM-att, and SENT-USD-Encoder-Decoder-LSTM. The weights metric $W$ was optimised using the Grey Wolf Optimization (GWO) metaheuristic algorithm.
4.3 GWO optimisation
The Grey Wolf Optimization (GWO) algorithm is employed at four different levels to optimise crucial aspects of the ensemble forecasting process: feature selection, sliding window size selection, model hyperparameter tuning, and final forecast blending. Generally, implementation of metaheuristic optimisations involves defining the main following procedures: problem preparation, objective function definition, population initialisation, and optimisation process. Firstly, the problem preparation involves specifying parameters to be calibrated and establishing their respective ranges (for instance, the range of sliding windows, the list of input features, and the number of epochs, see Table 1). The objective function representing the metric that the metaheuristic aims to either maximise or minimise during the optimisation process. This function typically encapsulates the performance of the deep learning model, such as accuracy or loss. The population in the context of metaheuristic optimisation refers to a collection of candidate solutions simultaneously. The metaheuristic explores and refines this population throughout the optimisation iterations to iteratively improve the model’s performance. A larger population generally provides more diversity and allows for a more thorough search space exploration, but it also increases computational cost.
@@ 4.3.1 Feature selection
The first parameter to be calibrated for each individual model using the GWO is the selection of input features. The range of possible values for this parameter includes four options: [None], [USDX], [SENT], and [USDX, SENT]. This allows the model to be evaluated using only historical Brent oil price data, USDX alone, SENT alone, or a combination of both features.
4.3.2 Sliding window size
This parameter represents the temporal scope of historical data used to train a model. Selecting an optimal sliding window size is a delicate balance between incorporating sufficient historical information for meaningful pattern recognition and avoiding overfitting or loss of relevance. Thus, selection suitable sliding window size is crucial to ensure the trained model learns from past data effectively. The defined range for calibrating this parameter spans from 3 to 30, allowing for exploration of the optimal value within this interval.
4.3.3 Neural networks hyperparameters tuning
Hyperparameter tuning is essential for achieving optimal performance in deep learning models, as it significantly affects training efficiency and accuracy. Finding the right hyperparameters, such as learning rate, optimizer, and hidden units, is challenging due to the complex interplay of parameters and the high-dimensional search space. We applied the GWO algorithm to find the optimal configurations for all the models described in Section 4.1. Table 1 presents the optimised hyperparameters and their corresponding range values.
@@ 4.3.4 Blending of final forecasts
Ensemble forecast methods can enhance accuracy by combining forecasts of diverse models that capture different aspects of the data. This often leads to better generalisation and robustness. However, the success of the ensemble methods depends on the individual models’ diversity and quality and the nature of the data. Thus, variant-Weight technique is a well-suited method for ensemble forecasting, in which multiple models are combined by assigning distinct weights to their respective predictions [26, 25]. To achieve this, we utilized the GWO algorithm to optimize the weight distribution, thereby determining the contribution of individual models to the final ensemble, as detailed in Section 4.2.
@@ 5 Experiments
5.1 Dataset and input features
The dataset used for the experiments consists of the daily observations of the Brent crude oil price for the period from (January 3, 2012, to April 1, 2021) including the period where the COVID-19 pandemic has affected energy and stock markets. It also includes Two external factors: the daily USD closing price (USDX) and a daily cumulative sentimental score (SENT).
Crude oil prices movement is heavily impacted by external factors that shape market dynamics. In our previous work [3], we analyzed the correlation between several external variables and Brent oil prices, including stocks, commodities, and sentiment analysis indicators. Our goal was to identify the most valuable variables for improving prediction accuracy while maintaining simplicity and efficiency. Accordingly, In this work, we integrate the U.S. dollar index (USDX) and sentimental index (SENT), individually and combined. The evident inverse relationship between Brent crude, which is traded in U.S. dollars, and the U.S. Dollar Index is a valuable factor in modeling Brent prices and is widely utilised in the literature [37]. While oil sentiment analysis shows substantial relevance to major events that impact the oil market. Thus, the sentimental index can reflect many other major factors such as supply and demand dynamics.
Historical data of Brent oil and USDX rate can be obtained from many public online sources such as Yahoo Finance. While SENT index is generated by the CrudeBERT model described in [38]. CrudeBERT is a variant of FinBERT that has been fine-tuned towards assessing the impact of market events on crude oil prices, focusing on frequently occurring market events and their impact on market prices according to Adam Smith’s theory of supply and demand. Mainly, CrudeBERT dataset used headlines originating from 1034 unique news sources, published on the Dow Jones newswires (approx. 21,200), Reuters (approx. 3,000), Bloomberg (approx. 1,100), and Platts (approx. 870). The sentimental score data is described and publicly available in [38]. The period from 2012 to 2021 was chosen for analysis due to the availability of sentiment scores during this time frame. Figure 3 displays the time-series plot of the daily closing price of Brent for this period.